Find the mean and standard deviation for each uniform continuous model. a. U(2,10) b. u(120, 200) c.u. (5, 94)
Answer #1

Ans! - Given continuous uniform and we have to find and standard deviation . disth out the mean lale know that Mean for contistandard deviation z Avernance 533 33 23:0940 c) uca=5 =g Pledn = (b+a)= 94+5 49.5 Vomance = (h-age = (94-53*= 792) 12 = 660-